# Annealing-based approach to solving partial differential equations

### v3

@@ Abstract
keywords:
Annealing-based approach , Partial differential equation , System of linear equations
[…]
[inst1]organization=Department of Computer Science, Ochanomizu University, addressline=2-1-1 Otsuka, Bunkyo-ku, city=Tokyo, postcode=112-8610, country=Japan
[…]
[inst2]organization=Graduate School of Information Sciences, Tohoku University, addressline=6-3-09 Aoba, Aramaki-aza Aoba-ku, city=Sendai, postcode=980-8579, state=Miyagi, country=Japan
1 Introduction
@@
2 Methods
2.1 Generalized eigenvalue problem
@@
[…]
@@
which has the same form as Eq. (2). Comparing Eq. (2) and Eq. (5), we identify $A=-\tilde{\bm{f}}\tilde{\bm{f}}^{\top}$ and $B=\tilde{K}$. Since $-\tilde{\bm{f}}\tilde{\bm{f}}^{\top}$ is a rank-$1$ matrix, Eq. (5) has only one nonzero eigenvalue, which is negative and corresponds to the smallest eigenvalue $\lambda_{\rm min}$, and the other eigenvalues are zero. Dividing both sides of Eq. (5) by $\tilde{\bm{f}}^{\top}\bm{v}_{\rm min}$, which is nonzero as $\lambda_{\rm min}\neq 0$, we obtain
@@
Considering that $\tilde{K}$ is positive definite, i.e., invertible, we arrive at the solution to Eq. (5):
@@ 2.2 Algorithm
$\bm{v}=\left[\begin{array}[]{c|c|c|c}-1,\frac{1}{2},\cdots,\frac{1}{2^{b-1}}&0&\cdots&0\\ \hline\cr 0&-1,\frac{1}{2},\cdots,\frac{1}{2^{b-1}}&&\vdots\\ \hline\cr\vdots&&\ddots&0\\ \hline\cr 0&\cdots&0&-1,\frac{1}{2},\cdots,\frac{1}{2^{b-1}}\end{array}\right]\bm{q}.$ (14)
Each element of $\bm{v}$ takes a discrete value in the range [1, $[-1,\;1-\frac{1}{2^{b-1}}]$.
@@
2.2.1 Initial guess stage
@@
2.2.2 Iterative descent stage
@@
[…]
@@
    Calculate the tentative eigenvector as $\bm{v}^{*}=\bm{v}+t\bm{d}$, where $t$ is defined by Eq. (23).
@@ 3 Results
where $x,y\in[0,1]$. The boundary conditions are $u(x,y)=0$ at $x=0,1$ or $y=0,1$. The exact solutions for these equations are $u(x)=x^{2}(x-1)^{2}$ for the one-dimensional symmetric case, $u(x)=x(x-1)(2x-1)(2x-3)/4$ for the one-dimensional asymmetric case, and $u(x,y)=xy(x-1)(y-1)$ for the two-dimensional case.
The system size of the SLE corresponding to the Poisson equation is determined by the dimension of vector $\bm{u}$ in Eq. (4). When the range $[0,1]$ is discretized into $n+1$ points in each dimension, the system size is $n$ and $n^{2}$ for one- and two-dimensional problems, respectively.
@@
3.1 Dependence on precision
Figure 1: The number of iterations required for solving the one-dimensional symmetric Poisson equation (Eq. (27)) as a function of the precision parameter $b$. (a) the initial guess stage. (b) iterative descent stage with $\eta=2^{-b+1}$. (c) iterative descent stage with $\eta=0.1$. The dimensions of solution vectors are $n=9$, $n=19$, and $n=39$. Filled symbols with solid lines correspond to short annealing times, and open symbols with dashed lines correspond to long annealing times. Symbols and shaded ranges represent the average and standard deviation, respectively. Symbol size is proportional to the success rate.
@@ 3.1 Dependence on precision
Figure 2: The number of precision updates required for solving the one-dimensional symmetric Poisson equation (Eq. (27)) as a function of the precision parameter $b$. (a) $b$-dependent mesh reduction rate ($\eta=2^{-b+1}$). (b) Constant mesh reduction rate ($\eta=0.1$). Black dashed lines indicate the theoretical estimate from Eq. (30) for $r=\epsilon_{0}=10^{-8}$. The dimensions of solution vectors are $n=9$, $n=19$, and $n=39$. Filled symbols with solid lines correspond to short annealing times, and open symbols with dashed lines correspond to long annealing times. Symbols and shaded ranges represent the average and standard deviation, respectively.
@@
Figure 3: The number of iterations at the iterative descent stage required for solving (a) the one-dimensional asymmetric Poisson equation (Eq. (28)) and (b) the two-dimensional Poisson equation (Eq. (29)) as a function of the precision parameter $b$. The dimensions of solution vectors are $n=9$, $n=19$, and $n=39$ for (a) and $n=3^{2}$, $n=5^{2}$, and $n=7^{2}$ for (b). Symbols and shaded ranges represent the average and standard deviation, respectively. Symbol size is proportional to the success rate.
@@
Comparing Fig. 3(a) to Fig. 1(b), we observe a significantly higher number of iterations, particularly for larger $n$. This difference can be attributed to the asymmetry of the exact solution $u(x)=x(x-1)(2x-1)(2x-3)/4$. Figure 3(a) indicates that solving more complex problems generally requires more iteration. Note that symbols for $b=4,8$ and $n=39$ are apparently small, which indicates low success rates.
In contrast, Fig. 3(b) shows a much smaller number of iterations for the two-dimensional Poisson equation, even for comparable system sizes $n$ to those in Fig. 1. This behavior can be attributed to the symmetric and convex nature of the exact solution $u(x,y)=xy(x-1)(y-1)$, which makes the optimization problem relatively easier. It is remarkable that the number of iterations decreases with increasing $b$ even for $n=7^{2}$.
3.2 Dependence on system size
[…]
@@ 5 Conclusions
Data availability
Data supporting the findings of this study are available upon request.
Acknowledgments
This work was supported by a JSPS KAKENHI Grant Number JP23H04499.

### v4

# Kazue Kudo

@@ Abstract
I Introduction
@@
II Methods
II-A Generalized eigenvalue problem
@@
where $\tilde{K}=K/\|\bm{f}\|$, and $\tilde{\bm{f}}=\bm{f}/\|\bm{f}\|$ is a unit vector. Equation (4) can be formulated as
@@
which has the same form as (2). Comparing (2) and (5), we identify $A=-\tilde{\bm{f}}\tilde{\bm{f}}^{\top}$ and $B=\tilde{K}$. Since $\tilde{\bm{f}}\tilde{\bm{f}}^{\top}$ is a rank-$1$ matrix, equation (5) has only one nonzero eigenvalue, which is negative and corresponds to the smallest eigenvalue $\lambda_{\rm min}$, and the other eigenvalues are zero. Dividing both sides of (5) by $\tilde{\bm{f}}^{\top}\bm{v}_{\rm min}$, which is nonzero as $\lambda_{\rm min}\neq 0$, we obtain
@@
Considering that $\tilde{K}$ is positive definite, i.e., invertible, we arrive at the solution to (5):
@@ II-B Algorithm
$\bm{v}=\left[\begin{array}[]{c|c|c|c}-1,\frac{1}{2},\cdots,\frac{1}{2^{b-1}}&0&\cdots&0\\ \hline\cr 0&\ddots&\ddots&\vdots\\ \hline\cr\vdots&\ddots&\ddots&0\\ \hline\cr 0&\cdots&0&-1,\frac{1}{2},\cdots,\frac{1}{2^{b-1}}\end{array}\right]\bm{q}.$ (14)
[…]
@@
II-B1 Initial guess stage
@@
II-B2 Iterative descent stage
@@
$t=\begin{cases}\mathrm{sgn}(b)\max\left(\frac{|b|}{a},\;1\right)&a>0,\\ 1&a\leq 0,\end{cases}$ (23)
@@
    Calculate the tentative eigenvector as $\bm{v}^{*}=\bm{v}+t\bm{d}$, where $t$ is defined by (23).
@@ III Results
where $x,y\in[0,1]$. The boundary conditions are $u(x,y)=0$ at $x=0,1$ or $y=0,1$. The exact solutions for these equations are $u(x)=x^{2}(x-1)^{2}$ for the one-dimensional symmetric case, u(x)=x(x1)(2x1)(2x3)/4u(x)=x(x-1)(2x-1)(2x-3)/4 for the one-dimensional asymmetric case, and u(x,y)=xy(x1)(y1)u(x,y)=xy(x-1)(y-1) for the two-dimensional case.
The system size of the SLE corresponding to the Poisson equation is determined by the dimension of vector $\bm{u}$ in (4). When the range $[0,1]$ is discretized into $n+1$ points in each dimension, the system size is $n$ and $n^{2}$ for one- and two-dimensional problems, respectively.
@@
III-A Dependence on precision
Figure 1: The number of iterations required for solving the one-dimensional symmetric Poisson equation (27) as a function of the precision parameter $b$. (a) the initial guess stage. (b) iterative descent stage with $\eta=2^{-b+1}$. (c) iterative descent stage with $\eta=0.1$. The dimensions of solution vectors are $n=9$, $n=19$, and $n=39$. Filled symbols with solid lines correspond to short annealing times, and open symbols with dashed lines correspond to long annealing times. Symbols and shaded ranges represent the average and standard deviation, respectively. Symbol size is proportional to the success rate.
@@ III-A Dependence on precision
Figure 2: The number of precision updates required for solving the one-dimensional symmetric Poisson equation (27) as a function of the precision parameter $b$. (a) $b$-dependent mesh reduction rate ($\eta=2^{-b+1}$). (b) Constant mesh reduction rate ($\eta=0.1$). Black dashed lines indicate the theoretical estimate from (30) for $r=\epsilon_{0}=10^{-8}$. The dimensions of solution vectors are $n=9$, $n=19$, and $n=39$. Filled symbols with solid lines correspond to short annealing times, and open symbols with dashed lines correspond to long annealing times. Symbols and shaded ranges represent the average and standard deviation, respectively.
@@
Figure 3: The number of iterations at the iterative descent stage required for solving (a) the one-dimensional asymmetric Poisson equation (28) and (b) the two-dimensional Poisson equation (29) as a function of the precision parameter $b$. The dimensions of solution vectors are $n=9$, $n=19$, and $n=39$ for (a) and $n=3^{2}$, $n=5^{2}$, and $n=7^{2}$ for (b). Symbols and shaded ranges represent the average and standard deviation, respectively. Symbol size is proportional to the success rate.
@@
Comparing Fig. 3(a) to Fig. 1(b), we observe a significantly higher number of iterations, particularly for larger $n$. This difference can be attributed to the asymmetry of the exact solution u(x)=x(x1)(2x1)(2x3)/4u(x)=x(x-1)(2x-1)(2x-3)/4. Figure 3(a) indicates that solving more complex problems generally requires more iteration. Note that symbols for $b=4,8$ and $n=39$ are apparently small, which indicates low success rates.
In contrast, Fig. 3(b) shows a much smaller number of iterations for the two-dimensional Poisson equation, even for comparable system sizes $n$ to those in Fig. 1. This behavior can be attributed to the symmetric and convex nature of the exact solution u(x,y)=xy(x1)(y1)u(x,y)=xy(x-1)(y-1), which makes the optimization problem relatively easier. It is remarkable that the number of iterations decreases with increasing $b$ even for $n=7^{2}$.
III-B Dependence on system size
Figure 4: The average number of iterations at the iterative descent stage for solving the one-dimensional symmetric, asymmetric, and two-dimensional Poisson equations as a function of system size $n$. Results are shown for $b=5$ and $N_{\rm step}=10^{3},\,10^{4}$. Symbol size is proportional to the success rate. The dashed line represents $ce^{\alpha n}$ with constants $c$ and $\alpha$.